OpenStudy (anonymous):

2NaN3 yields----> 3N2 + 2Na What volume of N2 gas measured at 1.3atm and 87C, would be produced by the reaction of 70g of NaN3?

OpenStudy (anonymous):

ok first you need to calculate amount of nitrogen gas in moles that can be created from 70 g of NaN3, firs calculate amount of NaN3 in moles then look at the ratio like this: n(NaN3) = m(NaN3)/M(NaN3) n(N2)/n(NaN3)= 3/2 n(N2)= 3/2 * m(NaN3)/M(NaN3) and then you just put it in pV=nRT (watch on units, they must be in kelvins and pascals) you should get that volume is 0,03635 m3 or 36,35 dm3